1) 1050 dollar is shared between A,B,C so that B receives 50% more than A, and C receives 25% more than B. How much does each receive?

2) A tank whose base is 2.25m square contains 9000 litres of water. Find the depth of the water.


1
Expert's answer
2021-10-20T14:21:25-0400

1) Let  A receives "x" dollar, then  B receives "1.5x" dollar, and  C receives "(1.5x)(1.25)" dollar. Then


"x+1.5x+(1.5x)(1.25)=1050"

"4.375x=1050"

"x=240"

"1.5x=360"

"(1.5x)(1.25)=450"

A receives "240" dollar, then  B receives "360" dollar, and  C receives "450" dollar.


2. Let "h=" the depth of the water.


"Base=2.25\ m^2"

"Volume=9000\ L=9\ m^3"

"Volume=Base(height)"

"height=h=\dfrac{Volume}{Base}=\dfrac{9\ m^3}{2.25\ m^2}=4\ m"

The depth of the water is 4 m.